# DUNIA: Pixel-Sized Embeddings via Cross-Modal Alignment for Earth Observation Applications

### v1

@@ Abstract
Significant efforts have been directed towards adapting self-supervised multimodal learning for Earth observation applications. However, existing methods produce coarse patch-sized embeddings, limiting their effectiveness and integration with other modalities like LiDAR. To close this gap, we present DUNIA, an approach to learn pixel-sized embeddings through cross-modal alignment between images and full-waveform LiDAR data. As the model is trained in a contrastive manner, the embeddings can be directly leveraged in the context of a variety of environmental monitoring tasks in a zero-shot setting. In our experiments, we demonstrate the effectiveness of the embeddings for seven such tasks (canopy height mapping, fractional canopy cover, land cover mapping, tree species identification, plant area index, crop type classification, and per-pixel waveform-based vertical structure mapping). The results show that the embeddings, along with zero-shot classifiers, often outperform specialized supervised models, even in low data regimes. In the fine-tuning setting, we show strong low-shot capabilities with performances near or better than state-of-the-art on five out of six tasks.
@@
Figure 1: DUNIA’s alignment strategy harnesses temporal and vertical information at the pixel level, enabling the model to develop a cross-modal understanding from a single satellite image. This approach serves as a foundation for cross-modal retrieval tasks across diverse Earth Observation (EO) applications. The pre-trained model is represented in navy blue, the multi-temporal autoencoder (AE) in yellow, and the waveform AE in green.
@@ 1 Introduction
With the rapid expansion of Earth Observation (EO) satellite missions, deep learning has emerged as a powerful solution for key EO applications, ranging from monitoring natural resources (Chen et al., 2023; Fayad et al., 2024; Li et al., 2023; Liu et al., 2022; Pauls et al., 2024; Schwartz et al., 2023) and assessing environmental impacts (Dalagnol et al., 2023; Wagner et al., 2023), to climate monitoring and forecast (Andrychowicz et al., 2023; Schultz et al., 2021), as well as agricultural and land resource management (Ienco et al., 2019; Kussul et al., 2017; Zhang et al., 2019).
While research of new deep learning architectures and upcoming remote satellite missions are expected to enhance existing EO applications, supervised deep learning approaches have key limitations: First, these models generally depend on labeled data, often requiring large annotated datasets. For example, the estimation of above-ground biomass (AGB) still heavily relies on traditional methods due to limited ground-based measurements and the need to use other variables as proxies (Morin et al., 2023; Schwartz et al., 2023).
@@
Foundation models (FMs) offer a promising solution to the aforementioned limitations by leveraging self-supervised pre-training on vast amounts of unlabeled data (Brown et al., 2020; Devlin, 2018; Kirillov et al., 2023; Minderer et al., 2022; Radford et al., 2021). For EO applications, such models are being developed to process diverse data types, including time series processing (Yuan et al., 2022), adaptation to different input satellite sensors (Xiong et al., 2024) or varying Ground Sample Distances (GSD) (Reed et al., 2023), and multi-modal data fusion (Astruc et al., 2025, 2024; Fuller et al., 2024). Current foundation models primarily address the adaptability issue, but not label scarcity. In fact, most EO-focused FMs are pre-trained masked auto-encoders (MAEs), which require extensive fine-tuning (Lehner et al., 2024; Singh et al., 2023). Moreover, existing FMs, even multi-modal ones, only leverage images. As such, they excel in EO tasks requiring horizontal structural understanding (e.g., land cover mapping, tree species identification), but struggle with EO applications requiring vertical structure understanding (e.g., canopy height estimation). Additionally, their patch-based approach limits direct pixel-level prediction capabilities. It also hinders the integration of LiDAR data, such as full-waveform LiDAR data available as labels at higher resolutions than the patch scale.
In this work, we propose a Dense Unsupervised Nature Interpretation Algorithm (DUNIA) that learns to generate pixel-level embeddings by aligning vertical structure information from a space-borne full waveform LiDAR with satellite imagery through contrastive learning. This dual alignment strategy enables both vertical and horizontal structure understanding, supporting diverse EO applications with minimal, if any, training. DUNIA achieves strong cross-modal retrieval performances, excels in forest structure benchmarks, and enables per-pixel vertical and horizontal structure retrieval. In our experiments, we demonstrate the effectiveness of the resulting embeddings for seven key EO applications and show that zero-shot classification operating on these embeddings can outperform specialized supervised EO models. A simplified version of our approach is presented in  Figure 1.
@@ 2 Background & Contribution
We sketch the related work along with the contributions made.
@@ 2.1 Contrastive Learning
Contrastive learning trains networks to create similar embeddings for related data and dissimilar embeddings for unrelated data, where related samples can be either different augmentations of the same input or inputs from different modalities. In this context, one of the main challenges is to avoid a so-called model collapse, in which case the model outputs become constant. This is typically addressed through (i) architectural designs or (ii) specialized objective functions.
Prominent examples of architectural designs are BYOL (Grill et al., 2020) and SimSiam (Chen & He, 2021), which respectively use asymmetry to prevent collapse, BYOL employing a momentum encoder and SimSiam using a stop-gradient operation with a predictor network. Specialized objective functions can be categorized into pair-based contrastive losses, clustering losses, and negative-pair-free losses. Here, pair-based losses, such as those presented in SimCLR (Chen et al., 2020) and MoCo (He et al., 2020), align positive pairs and separate negative pairs, but assume unique positive pairs within a mini-batch, which is not always true for EO data, especially at high resolutions. Clustering methods like SeLa (Asano et al., 2019) and SwAV (Caron et al., 2020) relax this constraint by performing the contrastive objective on data clusters. Negative-pair-free methods, such as Barlow Twins (Zbontar et al., 2021) and VICReg (Bardes et al., 2021), focus on redundancy reduction in the feature dimension. Zero-CL (Zhang et al., 2021) uses whitening transformations on the embeddings to maximize the trace of the cross-covariance matrix, achieving alignment and redundancy reduction without negative pairs.
@@ 2.2 Multi-Modal Learning
Multi-modal self-supervised approaches can be categorized into three main categories. (1) Modality-agnostic models use a shared network for different modalities (Carreira et al., 2022; Girdhar et al., 2022; Jaegle et al., 2021), but are typically limited to on one modality at a time during training, restricting cross-modal knowledge sharing (Srivastava & Sharma, 2024). For EO applications, several models follow this approach. Scale-MAE (Reed et al., 2023) and DOFA (Xiong et al., 2024) both process imagery across varying resolutions, the latter also processes different wavelengths using a common encoder. (2) Fusion-encoder models integrate data from multiple modalities through cross-modal attention, effectively combining information from each modality (Bachmann et al., 2022; Bao et al., 2021; Singh et al., 2022). OmniSat (Astruc et al., 2025) and AnySat (Astruc et al., 2024) exemplify this by integrating data from diverse EO sources into a unified representation. (3) Finally, multi-encoder approaches use separate encoders for each modality, which are then aligned in a shared latent space. This strategy is highly effective for tasks such as cross-modal retrieval and zero-shot classification across different modalities: image/text (Radford et al., 2021; Jia et al., 2021; Yuan et al., 2021; Yu et al., 2022), audio/text (Guzhov et al., 2022), and video/text (Luo et al., 2022; Pei et al., 2023).
@@ 2.3 Contribution
Our work aligns closely with the abovementioned third category, focusing on mono- and cross-modal retrieval within the multi-encoder approach. Unlike prior models that align paired data at the instance level (e. g., matching an image to its textual description), here we perform pixel-level alignment. This is necessary for dense predictions required for EO applications, and preserves the spatial resolution to match full LiDAR waveform footprints.
Specifically, our approach aligns sparse LiDAR waveforms with high-resolution imagery through contrastive learning to understand the vertical structure (i.e., pixel-waveform alignment) and simultaneously performs pixel-pixel alignment for horizontal structure understanding. DUNIA achieves strong cross-modal retrieval, surpasses specialized models in forest structure benchmarks, and enables per-pixel vertical structure retrieval. For land cover understanding, it performs competitively in zero-shot and low-shot evaluations. Additionally, its embeddings can directly generate waveforms representing the forest’s vertical structure from pixel inputs — a task not possible with existing methods.
@@ 3 Approach
Our goal is to learn two pixel-scale embedding spaces for two different EO data modalities using a single model, enabling simultaneous horizontal and vertical structural understanding. The vertical data come from GEDI (Dubayah et al., 2020), a spaceborne instrument that acquires sparse forest vertical structure measurements through 1D LiDAR waveforms and the horizontal data corresponds to 10 m resolution imagery from the Sentinel- 1 & 2 satellite missions (see Appendix B for additional details). Our framework is composed of two auto-encoders (AEs) and a third model that aligns, at the pixel-level, its output embeddings to the outputs of the two AEs. The two auto-encoders are, respectively, a multi-temporal image AE, and a full waveform AE. The third model is an encoder-decoder  type of model, with two similar, but separate decoders. Next, we provide some details on the main building blocks of the three main models, the pre-training objective and the latent diffusion model used for waveform generation. Full descriptions are deferred to Appendix A and the detailed model is presented in  Figure 2.
Figure 2: An overview of our framework showcasing the main blocks and connections. The pre-training model, which take as input as single date S-1 & 2 image mosaic, is shaded in blue. The multi-temporal image AE, which take as input a multi-date S-1& S2 image mosaic, is shaded in yellow. The vertical structure AE, which take as input the GEDI waveforms, is shaded in green.
3.1 Pre-training Model
The pre-training model in DUNIA (Figure 2) takes an image $\mathcal{I}\in\mathbb{R}^{H\times W\times C}$ and produces pixel level embeddings in $\mathbb{R}^{H\times W\times D_{p}}$. $C$ is the number of input channels, $(H,W)$ is the image resolution, and $D_{p}$ is the target projection dimension. $\mathcal{I}$ is a median composite image generated from Sentinel- 1 & 2 observations over several dates. The model comprises five building blocks: a patch embedding, a shared encoder, two similar decoders, neighborhood attention layers and several projection heads.
Following is a brief description of each of these blocks.
@@ 3.1 Pre-training Model
The 2D composite image $\mathcal{I}$ is split into $N=HW/P^{2}$ patches of dimension $N\times D$ using a convolutional layer, with $P$ and $D$ the patch size and the patch embedding dimension respectively.
@@
The encoder consists of 16 standard transformer layers (Vaswani et al., 2017), organized into 4 Transformer Blocks (TB), each containing four layers. Each layer takes and outputs a sequence of image tokens in $\mathbb{R}^{N\times D}$.
@@
The two decoders transform the encoder outputs into two sets of pixel-sized embeddings ($\mathcal{O}^{\mathcal{V}}$, $\mathcal{O}^{\mathcal{H}}\in\mathbb{R}^{H\times W\times D_{p}}$) for pixel-waveform and pixel-pixel alignment respectively. They use a hierarchical structure inspired by Ranftl et al. (2021) that progressively upsamples features while combining information from different encoder layers to maintain both fine and coarse details. Each decoder block ($d\in\{1,2,3,4\}$, with $d=1$ representing the final decoder layer) produces an image-like feature map of shape $\frac{H}{2^{(d-1)}}\times\frac{W}{2^{(d-1)}}\times D_{p_{d}}$, where $D_{p_{d}}=2^{p-1}D_{p}$ is the feature dimension out of block $d$. We refer to the output at these stages for a given decoder as $\mathcal{O}_{d}$. For simplicity, we will henceforth refer to $\mathcal{O}_{1}$ as $\mathcal{O}$.
@@
For the pixel-pixel alignment objective, we follow standard practice and append a projection head for each output of $\mathcal{O}^{\mathcal{H}}_{d}$. As the projection head becomes more specialized towards the training objective (Xue et al., 2024), its output becomes less generalizable when training and downstream objectives are misaligned, and as such it is discarded after training. For the pixel-waveform alignment, $\mathcal{O}^{\mathcal{V}}$ does not use a projection head, since the downstream tasks (e.g., waveform generation) align well with the training objective.
3.2 Waveform Processing Model
The waveform AE (Figure 2) is based on a VQ-VAE (Vector Quantized Variational Autoencoder) (van den Oord et al., 2017) architecture. It consists of three primary components: a waveform encoder, a residual vector quantizer (RVQ), and a waveform decoder. The input to this model are GEDI LiDAR waveforms ($\mathcal{W}$) available only for certain pixel locations in the composite image $\mathcal{I}$.

### v2

@@ Abstract
Significant efforts have been directed towards adapting self-supervised multimodal learning for Earth observation applications. However, most current methods produce coarse patch-sized embeddings, limiting their effectiveness and integration with other modalities like LiDAR. To close this gap, we present DUNIA, an approach to learn pixel-sized embeddings through cross-modal alignment between images and full-waveform LiDAR data. As the model is trained in a contrastive manner, the embeddings can be directly leveraged in the context of a variety of environmental monitoring tasks in a zero-shot setting. In our experiments, we demonstrate the effectiveness of the embeddings for seven such tasks: canopy height mapping, fractional canopy cover, land cover mapping, tree species identification, plant area index, crop type classification, and per-pixel waveform-based vertical structure mapping. The results show that the embeddings, along with zero-shot classifiers, often outperform specialized supervised models, even in low-data regimes. In the fine-tuning setting, we show strong performances near or better than the state-of-the-art on five out of six tasks.
@@
Figure 1: DUNIA’s alignment strategy leverages temporal and vertical information at the pixel level, enabling the model to develop a cross-modal understanding from a single satellite image. This approach serves as a foundation for cross-modal retrieval tasks across diverse EO applications. The pre-trained model is represented in navy blue, the multi-temporal autoencoder (AE) in yellow, and the waveform AE in green.
@@ 1 Introduction
With the rapid expansion of Earth Observation (EO) satellite missions, deep learning has emerged as a powerful solution for key EO applications, ranging from monitoring natural resources (Chen et al., 2023; Fayad et al., 2024; Li et al., 2023; Liu et al., 2022, 2023; Tolan et al., 2024; Lang et al., 2023; Pauls et al., 2024; Schwartz et al., 2023; Pauls et al., 2025) and assessing environmental impacts (Dalagnol et al., 2023; Wagner et al., 2023), to climate monitoring and forecast (Andrychowicz et al., 2023; Schultz et al., 2021), as well as agricultural and land resource management (Ienco et al., 2019; Kussul et al., 2017; Zhang et al., 2019).
While research of new deep learning architectures and upcoming remote satellite missions is expected to enhance existing EO applications, supervised deep learning approaches have key limitations. First, these models depend on labeled data, often requiring large annotated datasets. For example, the estimation of above-ground biomass (AGB) still heavily relies on traditional methods due to limited ground-based measurements and the need to use other variables as proxies (Morin et al., 2023; Schwartz et al., 2023).
@@
Foundation models (FMs) offer a promising solution to the aforementioned limitations by leveraging self-supervised pre-training on vast amounts of unlabeled data (Brown et al., 2020; Devlin, 2018; Kirillov et al., 2023; Minderer et al., 2022; Radford et al., 2021). For EO applications, such models are being developed to process diverse data types, including time series processing (Yuan et al., 2022), adaptation to different input satellite sensors (Xiong et al., 2024) or varying Ground Sample Distances (GSD) (Reed et al., 2023), and multi-modal data fusion (Astruc et al., 2025, 2024; Fuller et al., 2024). Most of current foundation models primarily address the adaptability issue, but not label scarcity. In fact, most EO-focused FMs are pre-trained masked auto-encoders (MAEs), which require extensive fine-tuning (Lehner et al., 2024; Singh et al., 2023). Moreover, existing FMs, even multi-modal ones, only leverage image-based modalities. As such, they excel in EO tasks requiring horizontal structural understanding (e.g., land cover mapping, tree species identification), but struggle with EO applications requiring vertical structure understanding (e.g., canopy height estimation). Additionally, their patch-based approach limits direct pixel-level prediction capabilities. This hinders the integration of LiDAR data, such as full-waveform LiDAR data available as labels at resolutions higher than the patch scale.
In this work, we propose a Dense Unsupervised Nature Interpretation Algorithm (DUNIA) that learns to generate pixel-level embeddings by aligning vertical (from a space-borne full waveform LiDAR) and horizontal structure information with satellite imagery through contrastive learning. This dual alignment strategy enables understanding of both vertical and horizontal structures, supporting diverse EO applications with minimal, if any, training. DUNIA achieves strong retrieval performance for several tasks and excels in forest structure benchmarks. In our experiments, we demonstrate the effectiveness of the resulting embeddings for seven key EO applications and show that zero-shot classification operating on these embeddings can outperform specialized supervised EO models in several cases. A simplified version of our approach is presented in  Figure 1.
@@ 2 Background & Contribution
We sketch the related works along with our contributions.
@@ 2.1 Contrastive Learning
Contrastive learning trains networks to produce similar embeddings for semantically related inputs and dissimilar embeddings for unrelated ones. Related inputs may include different augmentations of the same sample or paired examples from different modalities. In this context, one of the main challenges is to avoid model collapse, where model outputs become constant. This is typically addressed by: 1) architectural designs or 2) specialized objective functions.
Prominent examples of architectural designs are BYOL (Grill et al., 2020) and SimSiam (Chen & He, 2021), which respectively use asymmetry to prevent collapse, BYOL employing a momentum encoder and SimSiam using a stop-gradient operation with a predictor network. Specialized objective functions can be categorized into pair-based contrastive losses, clustering losses, and negative-pair-free losses. Here, pair-based losses, such as those presented in SimCLR (Chen et al., 2020) and MoCo (He et al., 2020), align positive pairs and separate negative pairs, but assume unique positive pairs within a mini-batch, which is not always true for EO data, especially at high resolutions. Clustering methods like SeLa (Asano et al., 2019) and SwAV (Caron et al., 2020) relax this constraint by optimizing a contrastive objective over cluster assignments instead. Negative-pair-free methods, such as Barlow Twins (Zbontar et al., 2021) and VICReg (Bardes et al., 2021), focus on redundancy reduction in the feature dimension. Zero-CL (Zhang et al., 2021) uses whitening transformations on the embeddings to maximize the trace of the cross-covariance matrix, achieving alignment and redundancy reduction without negative pairs.
2.2 Self-Supervised Learning for EO
Inspired by masked image modeling (MIM) from vision transformers, several SSL frameworks have adapted MAE-style objectives to the remote sensing domain. These methods mainly focus on reconstructing masked inputs to learn spatial, spectral, and temporal representations. SatMAE  (Cong et al., 2022) introduces a spectral-temporal aware masking scheme for Sentinel-2 (S-2) timeseries. SatMAE++  (Noman et al., 2024) extends SatMAE by incorporating multi-scale feature extraction. Scale-MAE  (Reed et al., 2023) embeds spatial resolution information into positional encodings for multiscale imagery. DOFA  (Xiong et al., 2024) extends masked autoencoding with a dynamic, spectral-aware weighting mechanism, enabling the model to encode diverse sensor modalities through a shared encoder.
Other related works on SSL for EO applications relies on a contrastive objective to capture the relationships within the data. CROMA  (Fuller et al., 2023) aligns radar and optical modalities using a cross-modal contrastive loss, supported by an auxiliary reconstruction head. DeCUR  (Wang et al., 2024) explicitly disentangles intra- and inter-modal contrastive signals based on Barlow Twins  (Zbontar et al., 2021). AnySat  (Astruc et al., 2024) proposed a resolution- and temporal-adaptive framework capable of handling data from various sensors and resolutions, effectively combining multiple EO data sources into a shared embedding space.
@@ 2.3 Multi-Modal Learning
Multi-modal SSL approaches can be categorized into three main categories. (1) Modality-agnostic models use a shared network for different modalities (Carreira et al., 2022; Girdhar et al., 2022; Jaegle et al., 2021), but are typically limited to one modality at a time during training, restricting cross-modal knowledge sharing (Srivastava & Sharma, 2024). For EO applications, several models follow this approach. Scale-MAE (Reed et al., 2023) and DOFA (Xiong et al., 2024) both process imagery across varying resolutions, the latter also processes different wavelengths using a common encoder. (2) Fusion-encoder models integrate data from multiple modalities through cross-modal attention, effectively combining information from each modality (Bachmann et al., 2022; Bao et al., 2021; Singh et al., 2022). OmniSat (Astruc et al., 2025) and AnySat (Astruc et al., 2024) exemplify this by integrating data from diverse EO sources into a unified representation. (3) Finally, multi-encoder approaches use separate encoders for each modality, which are then aligned in a shared latent space. This strategy is highly effective for tasks such as cross-modal retrieval and zero-shot classification across different modalities: image/text (Radford et al., 2021; Jia et al., 2021; Yuan et al., 2021; Yu et al., 2022), audio/text (Guzhov et al., 2022), and video/text (Luo et al., 2022; Pei et al., 2023).
@@ 2.4 Contribution
Our work aligns closely with the abovementioned third category, focusing on mono- and cross-modal retrieval within the multi-encoder approach. Unlike prior models that align paired data at the instance level (e. g., matching an image to its textual description), here we perform pixel-level alignment. This is necessary to preserve the spatial resolution required to match with full LiDAR waveform footprints.
Specifically, our approach aligns sparse LiDAR waveforms with high-resolution imagery through contrastive learning to understand the vertical structure (i.e., pixel-waveform alignment) and simultaneously performs pixel-pixel alignment for horizontal structure understanding. DUNIA enables per-pixel vertical and horizontal structure retrieval and surpasses specialized models in forest structure benchmarks. For land cover understanding, it performs competitively in zero-shot and low-shot evaluations. Additionally, pixel embeddings can directly generate waveforms representing forest vertical structures — a task not possible with existing methods.
@@ 3 Approach
Our objective is to learn two distinct pixel-level embedding spaces corresponding to two EO data modalities — vertical and horizontal — using a single pre-trained model. This enables simultaneous understanding of both horizontal and vertical structures. Vertical information is derived from the Global Ecosystem Dynamics Investigation (GEDI) instrument (Dubayah et al., 2020), a spaceborne LiDAR that provides sparse measurements of 3D structures via 1D waveform signals. Horizontal information comes from 10 m resolution imagery acquired by the S-1 & 2 missions.
While our framework can be implemented in various ways, we adopt a modular design centered around a transformer encoder and two independent convolutional decoders. This encoder-decoder architecture enables us to generate pixel-sized embeddings for each input. The two decoders are designed to disentangle horizontal and vertical structure understanding: one focuses on spatial (horizontal) relationships, the other on vertical structure.
For horizontal structure alignment, we use a multi-temporal autoencoder (AE) that encodes a sequence of satellite images and produces temporally informed, per-pixel embeddings. This design achieves two goals: 1) it provides an augmented input to the pre-trained model, and 2) it allows the model to implicitly capture phenological patterns without requiring a time series at inference time.
For vertical structure alignment, we also use an AE, this time encoding GEDI waveforms. The waveform encoder projects the 1D input into a latent space, which is aligned with the corresponding pixel embedding from the pre-trained model. While the decoder is not needed for the alignment itself, it is necessary for waveform generation, ensuring that the latent space remains semantically and structurally meaningful.
Figure 2 illustrates the main components of our approach, including the inputs to each model, their architectural building blocks, and the layers used for alignment. Below, we describe the main modules in more detail, followed by our pre-training objectives and the latent diffusion model for waveform generation. Full details are provided in Appendix A.
Figure 2: Simplified architectural overview of the proposed framework. The figure illustrates the key encoder and decoder components, input modalities, and the layers where alignment is performed (red dashed arrows). The pre-trained model is shaded in blue. The multi-temporal image AE, is shaded in yellow. The waveform encoder is shaded in green. Additional components, losses and training objectives are omitted for clarity; see Figure A1 for full details.
3.1 Pre-trained Model
The pre-trained model in DUNIA (Figure 2) takes an image $\mathcal{I}\in\mathbb{R}^{H\times W\times C}$ and produces pixel level embeddings in $\mathbb{R}^{H\times W\times D_{p}}$. $C$ is the number of input channels, $(H,W)$ is the image resolution, and $D_{p}$ is the target projection dimension. $\mathcal{I}$ is a median composite image generated from S-1 & 2 observations over several dates. The pre-trained model comprises five building blocks: a patch embedding, a shared encoder, two similar decoders, neighborhood attention layers and several projection heads. Following is a brief description of each of these blocks.
@@ 3.1 Pre-trained Model
The 2D composite image $\mathcal{I}$ is split into $N=HW/P^{2}$ patches, each of dimension $D$ using a convolutional layer, with $P$ and $D$ the patch size and the patch embedding dimension respectively.
@@
The encoder consists of 16 standard transformer layers (Vaswani et al., 2017), organized into four Transformer Blocks, of four layers each. Each layer inputs/outputs a sequence of image tokens in $\mathbb{R}^{N\times D}$.
@@
The two decoders transform the encoder outputs into two sets of pixel-sized embeddings ($\mathcal{O}^{\mathcal{V}}$, $\mathcal{O}^{\mathcal{H}}\in\mathbb{R}^{H\times W\times D_{p}}$) for pixel-waveform and pixel-pixel alignment respectively. They use a hierarchical structure inspired by Ranftl et al. (2021) that progressively upsamples features while combining information from different encoder layers to maintain both fine and coarse details. Each decoder block ($d\in\{1,2,3,4\}$, with $d=1$ representing the final decoder layer) produces an image-like feature map of shape $\frac{H}{2^{(d-1)}}\times\frac{W}{2^{(d-1)}}\times D_{p_{d}}$, where $D_{p_{d}}=2^{d-1}D_{p}$ is the feature dimension out of block $d$. We refer to the output at these stages for a given decoder as $\mathcal{O}_{d}$. For simplicity, we will henceforth refer to $\mathcal{O}_{1}$ as $\mathcal{O}$.
@@
For the pixel-pixel alignment objective, we follow standard practice and append a projection head for each output of $\mathcal{O}^{\mathcal{H}}_{d}$. As the projection head becomes more specialized towards the training objective (Xue et al., 2024), its output becomes less generalizable when training and downstream objectives are misaligned, and as such it is discarded after training. For the pixel-waveform alignment $\mathcal{O}^{\mathcal{V}}$ no projection head is used as the downstream tasks (e.g., waveform generation) align with the training objective.
3.2 Waveform Model
The waveform AE is based on a VQ-VAE (Vector Quantized Variational Autoencoder) (van den Oord et al., 2017) architecture. It consists of three primary components: a waveform encoder (Figure 2), a residual vector quantizer (RVQ), and a waveform decoder. The input to this model are GEDI LiDAR waveforms ($\mathcal{W}$) available only for certain pixel locations in the composite image $\mathcal{I}$.